OpenStudy (anonymous):

medal and much blessings and thank you If x varies inversely with y and x = 4 when y = 9, find x when y = 6. 8 6 3 i got 6

OpenStudy (anonymous):

y varies directly with x y=kx 10=k(15) k=10/15 y=(10/15)x when x=6 y=(10/15)(6)=60/15=4 x varies inversely as y x=k/y 2=k/8 k=16 x=16/y when y=17 x=16/17

OpenStudy (ikram002p):

when its directly simply use this :-